A department uses the weighted average method of process costing. All direct materials are added at the beginning of the process. This department has the following data for this month.

  • 189 units in beginning WIP (30% complete with respect to conversion costs).
  • 10,323 units started
  • 214 units in ending WIP (46% complete with respect to conversion costs).
  • $2,145 beginning WIP direct materials costs
  • $1,493 beginning WIP conversion costs
  • $85,029 current month direct materials costs
  • $85,642 current month conversion costs

What is the department's conversion cost per equivalent unit for this month (round final answer to nearest cent if necessary)?
Answer #1

Unit transferred out = 189+10323-214 = 10298

Equivalent unit of conversion = 10298+(214*46%) = 10396.44

Total Cost of conversion = (1493+85642) = 87135

Cost per equivalent unit of conversion = 87135/10396.44 = 8.38